OpenStudy (anonymous):

what is the common difference for the geometric sequence 4, 12, 36

OpenStudy (anonymous):

I really need help here

OpenStudy (ash2326):

A geometric sequence is of the form \[a, ar, ar^2...ar^n\] Two consecutive terms are in the same ratio that is \[\frac{ar}{a}=\frac{ar^2}{ar}=\frac{ar^3}{ar^2}=r\] here r is the common difference or ratio so here we have 4, 12 and 36 so \[\frac{12}{4}=\frac{36}{12}\] Can you find this?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

okay thank you so much

OpenStudy (ash2326):

do you understand?

OpenStudy (anonymous):

yeah no i get it it 3*4=12*3=36 and so on

OpenStudy (ash2326):

good

OpenStudy (anonymous):

thanks